OpenStudy (anonymous):

logbase4(0.25)+logbase2(0.0625)

OpenStudy (tkhunny):

Really Big Hints \(0.25 = 4^{-1}\) \(0.0625 = 2^{-4}\)

OpenStudy (anonymous):

please tell me the final answer

OpenStudy (tkhunny):

Please read the Code of Conduct. There should be a link on the lower left of your screen. I am constrained from simply giving answers. Show your work and let's talk.

OpenStudy (anonymous):

\[\log_{4} 0.25+\log_{2} 0.0625 is\]

OpenStudy (tkhunny):

Retyping the problem statement does not constitute any work shown. A logarithm IS an exponent. Those were REALLY BIG hints. Give it some thought.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my answer is 5/16 is that right

OpenStudy (tkhunny):

Writing down an answer does not constitute showing work. No, 5/16 is not correct. Show my how you managed that answer.

OpenStudy (anonymous):

\[4^{-1}+2^{-4}=1/4+1/2^{4}=1/4+1/16=5/16\]

OpenStudy (tkhunny):

Why did you just discard the logs? There is no rule like that. You must use rules. \(\log_{4}(0.25) = \log_{4}\left(4^{-1}\right) = (-1)\cdot\log_{4}(4) = (-1)\cdot (1) = -1\) There is no shortcut. You do the other piece.

OpenStudy (anonymous):

ok I got that but I hv a question hoe did u get \[\log_{4}0.25=\log_{4}(4^{-1}) \]

OpenStudy (tkhunny):

\(0.25 = \dfrac{1}{4} = 4^{-1}\)

OpenStudy (anonymous):

\[\log_{2}0.0625=\log_{2}(0.25)^{2}=2\log_{2}0.25=2\log_{2}(1/4)=2\log_{2}(4^{-1})=2\log_{2}(2^{-2})=2x-2\log_{2}^{2}=-4x1=-4 \]

OpenStudy (tkhunny):

Ran off the edge. That's fine but it's a little laborious. Since the Base is 2, I am caused to wonder if there is an integer power of 2 in that argument. \(2^{-1} = 0.5\) \(2^{-2} = 0.25\) \(2^{-3} = 0.125\) \(2^{-4} = 0.0625\) -- Aha!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

another question \[\log_{\sqrt{5}}0.008+\log_{2*\sqrt[3]{5?}} 1600 1s \]

OpenStudy (anonymous):

\[\log_{\sqrt{5}} 0.008\]

OpenStudy (tkhunny):

Wow! What's the point of this question? Anyway, I might be tempted to transform to a common base. \(\log_{b}(a) = \dfrac{\log_{c}(a)}{\log_{c}(b)}\) for suitable a, b, and c.

OpenStudy (tkhunny):

Or not. \(0.008 = 8/1000 = 4/500 = 2/250 = 1/125 = 5^{-3}\)

OpenStudy (tkhunny):

Do you see it?

OpenStudy (anonymous):

okay I got that

OpenStudy (tkhunny):

Then \(5 = (\sqrt{{5}})^{2}\), that \(5^{-3} = (\sqrt{5})^{-6}\) and we're almost done. I need to go. You tackle that other piece. Keep in mind, since it has been the nature of all the problems we have seen, that you may just want to calculate some integer exponents of the base and see how long it takes to find the argument.

OpenStudy (anonymous):

1. \[\log_{8}128-\log_{9}\cot (\pi/3) \] is